# LLM-Driven Corrective Robot Operation Code Generation with Static Text-Based Simulation

### v1

@@ III-C Static Text-Based Simulation using LLM
Given the robot operation code generated by the code generator, the goal of our static text-based simulation is to accurately interpret the robot’s actions, reason about the state transitions, predict the robot states, and generate an observation of the robot’s trajectory. Specifically, we formulate our text-based simulation using LLM as $O=\langle S,A,T,C\rangle$, where $O$ denotes the robot trajectory observation produced by the simulation, $S=(s_{0},s_{1},s_{2},\dots,s_{n})$ denotes the finite set of discrete states, AA denotes the finite set of robot actions, $T:S\times A\rightarrow S$ denotes the state transition, and $C$ denotes the text of robot operation code. The trajectory observation $O$ captures a sequence of robot actions $l=(a_{1},a_{2},\dots,a_{n})$ that transitions robot from the initial state $s_{0}$ through intermediate states, to the final state $s_{n}$. $O$ also embeds the history of the robot’s trajectory dynamics for each robot action $l$ and their transitions $T$ after executing code $C$. This observation enables the subsequent evaluation to identify the mismatches and correct them in the next code generation. Therefore, the reliability of LLM-driven robotics depends critically on the accuracy of $O$.
@@ IV-C2 Text-Based Simulation Accuracy
To further evaluate the reliability of our text-based simulation, we compare the accuracy of the trajectory observations generated by the LLM-simulator against those obtained using the Semantic method [CLGSCE] in the dynamic simulator. For each task in the Advanced task set, we construct a corresponding ground truth code ($C^{\text{correct}}$) that implements the task. The LLM-simulator is then used to statically simulate the execution of each $C^{\text{correct}}_{i}$ and produce trajectory observations. The accuracy of the simulation is computed as $TP/N\times 100\%$, where TPTP denotes the number of observations that accurately capture the robot trajectory dynamics, and NN is the total number of simulated code in $C^{\text{correct}}$.
@@ IV-C3 Text-Based Simulation Evaluation Accuracy
We further analyze whether the observations generated by the LLM-simulator can effectively support the evaluation process. Specifically, we compare the evaluation accuracy when using observations generated by our LLM-simulator against those produced from the Semantic method [CLGSCE]. This experiment leverages the ground-truth code set $C^{\text{correct}}$ in section IV-C2, along with an additional set $C^{\text{incorrect}}$ that contains at least one error action in code for each task. The evaluation accuracy is computed as $(TP+TN)/N\times 100\%$, where TPTP denotes the number of correct evaluations on $C^{\text{correct}}$ and $C^{\text{incorrect}}$ (i.e., correctly identifying whether the trajectory matches with the task and providing a detailed explanation if mismatches are identified), and NN is the total number of evaluated code scripts.

### v2

@@ III-C Static Text-Based Simulation using LLM
Given the robot operation code generated by the code generator, the goal of our static text-based simulation is to accurately interpret the robot’s actions, reason about the state transitions, predict the robot states, and generate an observation of the robot’s trajectory. Specifically, we formulate our text-based simulation using LLM as $O=\langle S,A,T,C\rangle$, where $O$ denotes the robot trajectory observation produced by the simulation, $S=(s_{0},s_{1},s_{2},\dots,s_{n})$ denotes the finite set of discrete states, $A$ denotes the finite set of robot actions, $T:S\times A\rightarrow S$ denotes the state transition, and $C$ denotes the text of robot operation code. The trajectory observation $O$ captures a sequence of robot actions $l=(a_{1},a_{2},\dots,a_{n})$ that transitions robot from the initial state $s_{0}$ through intermediate states, to the final state $s_{n}$. $O$ also embeds the history of the robot’s trajectory dynamics for each robot action $l$ and their transitions $T$ after executing code $C$. This observation enables the subsequent evaluation to identify the mismatches and correct them in the next code generation. Therefore, the reliability of LLM-driven robotics depends critically on the accuracy of $O$.
@@ IV-C2 Text-Based Simulation Accuracy
To further evaluate the reliability of our text-based simulation, we compare the accuracy of the trajectory observations generated by the LLM-simulator against those obtained using the Semantic method [CLGSCE] in the dynamic simulator. For each task in the Advanced task set, we construct a corresponding ground truth code ($C^{\text{correct}}$) that implements the task. The LLM-simulator is then used to statically simulate the execution of each $C^{\text{correct}}_{i}$ and produce trajectory observations. The accuracy of the simulation is computed as $TP/N\times 100\%$, where $TP$ denotes the number of observations that accurately capture the robot trajectory dynamics, and $N$ is the total number of simulated code in $C^{\text{correct}}$.
@@ IV-C3 Text-Based Simulation Evaluation Accuracy
We further analyze whether the observations generated by the LLM-simulator can effectively support the evaluation process. Specifically, we compare the evaluation accuracy when using observations generated by our LLM-simulator against those produced from the Semantic method [CLGSCE]. This experiment leverages the ground-truth code set $C^{\text{correct}}$ in section IV-C2, along with an additional set $C^{\text{incorrect}}$ that contains at least one error action in code for each task. The evaluation accuracy is computed as $(TP+TN)/N\times 100\%$, where $TP$ denotes the number of correct evaluations on $C^{\text{correct}}$ and $C^{\text{incorrect}}$ (i.e., correctly identifying whether the trajectory matches with the task and providing a detailed explanation if mismatches are identified), and $N$ is the total number of evaluated code scripts.